OpenStudy (anonymous):

Jill arranges quarters in stacks of increasing powers of 2, starting with the zeroth power (2^0). If she ends up with six stacks, how many coins has she used ? HELP ME PLEASE !!!

OpenStudy (wolf1728):

1st stack = 2^0 =1 2nd = 2^1 = 2 3rd = 4 4th = 8 5th = 16 6th = 32 Total all six = 63 quarters

OpenStudy (anonymous):

Thanks